# Hierarchical LLMs In-the-loop Optimization for Real-time Multi-Robot Target Tracking under Unknown Hazards

### v1

@@ IV-B2 Modular Evaluation
We conducted a comparison of three approaches: one without any LLMs (w/o LLMs), another with LLMs in the loop (w/ LLMs), and one with both LLMs and human input into the system. We consider a scenario with two drones tracking four targets in the same direction. We specify targets to have different velocities to increase the complexity of the conditions. The environment includes two sensing danger zones and one communication danger zone. The human input is included at a lower frequency to push the LLMs to improve the performance, the prompt is
@@
Focus more on tracking targets; The trace is not good. We utilized GPT-4o for better performance, with the results presented in Tab. I. By adding the LLMs in the loop to adjust the parameters, the system can achieve better tracking performance and get fewer attacks from sensing and communication danger zones. With human input to express concern about the performance, LLMs can dramatically decrease the tracking errors computed by the trace of the estimation covariance matrix.
@@
IV-C Hardware Experiments
We validated the ability of the proposed framework through experiments with a team of four robots in both indoor and outdoor scenarios, as shown in Fig. 1 and 7. The outdoor experiments include more challenges like noisy detection, estimation, and control under disturbance. In particular, we set up two Scarab ground robots [38] to serve as moving targets, with each carrying an AprilTag [39] on top. To effectively detect and track the moving targets, two Dragonfly 230 MAVs [40] were employed. The Dragonfly 230 MAV is equipped with a downward-facing VGA camera for state estimation and target detection. An Unscented Kalman Filter is running on board the MAV to fuse the 30Hz VIO odometry from the VOXL board with the IMU to get 150Hz pose information. We used an additional AprilTag to create a common reference frame for all MAVs. Once an aerial agent observes the tag, the agent estimates the relative transformation between the common reference frame and its odometry frame and registers the detected target in the common reference frame. The robot receives global waypoints from the ground station and transforms them into its odometry frame for trajectory generation and control. The state estimation, tag detection, trajectory planning, and controller run onboard the drones.
(a) Indoor
(b) Outdoor
@@ IV-C Hardware Experiments
To address unexpected changes and concerns regarding environmental conditions or hazards, the following examples are provided:
@@
The human supervisor has inputs: The robots seem to not track the targets.
In response, the Task LLM adjusted the assignment, refocusing each drone on a single target. The updated plan was as follows:
@@ V Conclusion
In this work, we propose a hierarchical framework that combines the reasoning capabilities of LLMs with traditional optimization techniques for multi-robot target tracking. Our results demonstrate the capacity of this framework to handle unexpected conditions, showing improved performance in resilience, adaptability, and tracking accuracy, particularly in dynamic and hazardous environments. This highlights the potential of LLM-integrated robotics systems in a wide range of real-world applications.

### v2

@@ IV-A3 Ablation Study
We compare three approaches: a baseline system using expert-defined parameters and fixed task allocation without any LLM (w/o LLMs), one with LLMs in the loop (w/ LLMs), and one with LLMs and human input into the system (w/ LLMs + Human). We consider a scenario with two drones tracking four targets in the same direction. We specify targets to have different velocities to increase the complexity of the conditions. The environment includes two sensing danger zones and one communication danger zone. Human input is provided at a lower frequency to encourage the LLMs to independently improve their performance through adaptation. An example prompt is:
Focus more on tracking targets; The trace is not good.
@@
To illustrate, we use GPT-4o with the results presented in Tab. I. By adding the LLMs in the loop to adjust the parameters, the system can achieve better tracking performance and get fewer attacks from sensing and communication danger zones (H3). With human input to express concern about the performance, LLMs can dramatically reduce tracking errors computed by the trace of the estimation covariance matrix.
@@
IV-B Hardware Experiments
We validated the ability of the proposed framework through experiments with a team of four robots in both indoor and outdoor scenarios (H4), as shown in Fig. 1 and 5. The outdoor experiments include more challenges like noisy detection, estimation, and control under disturbance. We set up two ground robots to serve as moving targets, with each carrying an AprilTag [41] on top. In hardware experiments, the update frequency of each module was adjusted based on the specific scenario and runtime constraints. For example, in a real-world deployment involving two aerial drones tracking two ground robots, the computation time and update frequency of each major module are summarized in Tab. II. The action LLM runs 4–5 times more frequently than the task LLM, enabling more responsive low-level adaptations. To support long-horizon reasoning, the task LLM received a history of robot and environment states from the past 5 steps. If the LLM output is invalid or violates constraints, the system retries with adjusted prompts until a valid response is obtained, ensuring robustness to occasional failures.
To effectively detect and track the moving targets, two drones were employed. The drone is equipped with a downward-facing VGA camera for state estimation and target detection. An Unscented Kalman Filter is running on board the drone to fuse the 30Hz VIO odometry with the IMU to get 150Hz pose information. We used an additional AprilTag to create a common reference frame for all drones. Once an aerial agent observes the tag, the agent estimates the relative transformation between the common reference frame and its odometry frame and registers the detected target in the common reference frame. The robot receives global waypoints from the ground station and transforms them into its odometry frame for trajectory generation and control. The state estimation, tag detection, trajectory planning, and controller run onboard the drones.
TABLE II: Module Execution During a Single Experiment Trial
Module Frequency (Hz) Avg. Runtime (s) Success Rate
Task LLM 0.05 1.3639 100%
Action LLM 0.2 0.8911 100%
Optimizer 10 0.0034 100%
@@ IV-B Hardware Experiments
To address unexpected changes regarding environmental conditions or hazards, the following examples are provided:
@@
The human supervisor has inputs: The robots do not seem to track the targets.
In response, the Task LLM reassigned and refocused each drone to a single target. The updated plan was:
@@ V Conclusion
In this work, we propose a hierarchical framework that combines the reasoning capabilities of LLMs with optimization techniques for multi-robot target tracking. Our results demonstrate the capacity of this framework to handle unexpected conditions, showing improved performance in resilience, adaptability, and tracking accuracy, particularly in dynamic and hazardous environments. This highlights the potential of LLM-integrated robotics systems in a wide range of real-world applications.
This work explores a promising direction for optimization-informed LLMs, but also presents limitations. LLM inference introduces latency and relies on carefully crafted prompts, structured optimization formulations, and constrained outputs, which pose challenges for scaling to larger or more reactive multi-robot systems. Future work will focus on improving scalability and robustness while maintaining lightweight onboard deployment capabilities.